OpenStudy (mikeymaximum):

Use ΔDEF to answer the question. What is tan D? A. 8/17 B. 8/15 C. 15/17 D. 15/8

Nnesha (nnesha):

\[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

 which one is opposite side of angle D ?